OpenStudy (andu1854):

A triangle has a side of c=2 and angles A = pi/4 and B = pi/3. Find the length of a of the side opposite A.

OpenStudy (anonymous):



OpenStudy (anonymous):

we can solve angle C, then use sine law to find a.

OpenStudy (anonymous):

sum of all angles in a triangle is equal to pi radian or 180 deg.

OpenStudy (anonymous):

@andu1854 do you know how to do it?

OpenStudy (souvik):

\[\frac{ \sin A }{ a } =\frac{ \sin C }{ c }\] you know the angle C...you can find the value of a